OpenStudy (anonymous):

How far apart are the foci of an ellipse with a major axis of 26 feet and a minor axis of 24 feet? 4 feet 5 feet 10 feet 20 feet

OpenStudy (mertsj):

Did you find c?

OpenStudy (anonymous):

Not yet, is this right so far? 2a=26 and 2b=24

OpenStudy (mertsj):

yes

OpenStudy (anonymous):

Then I do a^2+b^2=c^2?

OpenStudy (anonymous):

When a=13 and b=12

OpenStudy (mertsj):

I believe that is for a hyperbola. This is an ellipse.

OpenStudy (anonymous):

a^2-b^2=c^2?

OpenStudy (mertsj):

yes

OpenStudy (anonymous):

c=5

OpenStudy (anonymous):

is that my answer?

OpenStudy (mertsj):

yes

OpenStudy (anonymous):

thanks!

OpenStudy (mertsj):

So how far apart are the foci?

OpenStudy (anonymous):

5 Feet

OpenStudy (mertsj):

tsk tsk

OpenStudy (mertsj):

One is (c,0) and the other is (-c,0)

OpenStudy (anonymous):

or wait is it 10

OpenStudy (mertsj):

That's better.